OpenStudy (anonymous):

A watermelon composed of 99% of water, weighted 10kg. After a while, some of water evaporated, leaving the watermelon composed of 98% of water. What is the new weight of water-melon?

OpenStudy (anonymous):

9kg?

OpenStudy (anonymous):

99/10 = 98/x (99% to 10 kg = 98% to x kg) cross multiply 99x = 980 x = 980/99 x = 9.89 = 9.9 kg I am not 100% sure on this...you will need a second opinion

OpenStudy (dumbcow):

i concur